OpenStudy (john123):

y = (x + 2)^2 + 4 in standard form ?

OpenStudy (anonymous):

y = (x + 2)(x + 2) + 4 FOIL THIS y = x^2 + 2x + 2x + 4 + 4 ADD LIKE TERMS y = x^2 + 4x + 8